### v1

@@ 3 Efficient Douglas-Rachford Algorithms on Hadamard Manifolds
Initialization: Choose $x_{0}$ arbitrary.
@@
$\begin{cases}y_{n}=\gamma(x_{n},{R}_{\lambda\phi}{R}_{\lambda\psi}(x_{n});\alpha_{n}),\\ x_{n+1}=({R}_{\lambda\phi}{R}_{\lambda\psi})^{p}(y_{n})\text{ for all }n\in\mathbb{N}_{0},\end{cases}$ (45)
@@
Since $T={R}_{\lambda\phi}{R}_{\lambda\psi}$ is nonexpansive. From Theorem 22 sequence $\{x_{n}\}$ generated by (44) converges to an element $v\in\operatorname*{Fix}({R}_{\lambda\phi}{R}_{\lambda\psi})$. From Theorem 16, $u=\operatorname*{Prox}_{\lambda\psi}(v)$ is solution of problem (2). ∎
@@
Since $T={R}_{\lambda\phi}{R}_{\lambda\psi}$ is nonexpansive. From Theorem 27 sequence $\{x_{n}\}$ generated by (45) converges to an element $v\in\operatorname*{Fix}({R}_{\lambda\phi}{R}_{\lambda\psi})$. From Theorem 16, $u=\operatorname*{Prox}_{\lambda\psi}(v)$ is solution of problem (2). ∎
@@ 4 Numerical Experiments
where $a,b\in(0,\infty)$ (see [52]). In order to handle the nonconvex problem (46) in $\mathcal{M}=\mathbb{R}^{2}$, consider
@@
Then $\mathcal{M}$ is a Hadamard manifold with respect to metric (47) (see [53]) and the following statements hold:
@@
Thus $\phi$ and $\psi$ are geodesically convex on $\mathcal{M}$. For more details of this splitting; see [54].
@@
    For $\lambda>0$, the proximal mapping $\operatorname*{Prox}_{\lambda\phi}$ and $\operatorname*{Prox}_{\lambda\psi}$ are defined by
    $\operatorname*{Prox}_{\lambda\phi}(x)=\begin{pmatrix}x_{1}\\ \frac{x_{2}+2a\lambda x_{1}^{2}}{1+2a\lambda}\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}$
    and
    $\operatorname*{Prox}_{\lambda\psi}(x)=\begin{pmatrix}\frac{x_{1}+2\lambda b}{1+2\lambda}\\ x_{2}-\frac{4\lambda(x_{1}+2\lambda b)(x_{1}-b)+4\lambda^{2}(x_{1}-b)^{2}}{(1+2\lambda)^{2}}\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}.$
@@
    $R_{\lambda\phi}(x)=\begin{pmatrix}x_{1}\\ \frac{2(x_{2}+2a\lambda x_{1}^{2})}{1+2a\lambda}-x_{2}\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}$ (49)
    and
@@
Remark 35.
The reflection operators $R_{\lambda\phi}$ and $R_{\lambda\psi}$ defined by (49) and (50), respectively are nonexpansive; see Appendix B.
@@
$\min_{x\in\mathcal{M}}\phi(x)+\psi(x).$ (51)
For the numerical experiment, we compare three different methods: the Douglas-Rachford method [Alg(DR)] (3), the inertial Douglas-Rachford method [Alg(InDR)] (44) and the $p$-accelerated normal S-Douglas-Rachford method [Alg($p$-AccDR)] (45). All numerical tests are carried out using MATLAB R2024b on an Apple Macbook Air Laptop with a 1.8 GHz Dual-Core Intel Core i5 processor and 8 GB RAM.
We set $a=1$, $b=2$ and $\lambda=1$. We take the initial point $x_{0}=(1,2)^{T}$ for all algorithms and $x_{1}=(1,3)^{T}$ for Alg(InDR). We choose the iteration parameters $\alpha_{n}=0.5$ and $\theta_{n}=0.3$. All algorithms stops when they meet the stopping criteria $E_{n}=d(\operatorname*{Prox}_{\lambda\psi}(x_{n}),\operatorname*{Prox}_{\lambda\psi}(x_{n-1}))<10^{-14}$.
Parameters for these three algorithms are listed below:
[Alg(DR)] $:$ $\alpha_{n}=0.5;$
[Alg(InDR)] $:$ $\alpha_{n}=0.5,\theta_{n}=0.3;$
[Alg($p$-AccDR)] $:$ $\alpha_{n}=0.5,\beta_{n}=0.3.$
The runtime and the number of iterations are depicted in Table 1 and the graph of values of $E_{n}$ with respect to the number of iterations $n$ is shown in Figure 2. From Table 1 and Figure 2, we can see that Alg($p$-AccDR) is faster than [Alg(InDR)] and [Alg(DR)] in both runtime and number of iterations.
Algorithm Runtime(sec) Iterations
Alg(DR) 0.083934 67
Alg(InDR) 0.034095 32
Alg(pp-AccDR) 0.011514 16
Table 1: Numerical performance of the runtime and number of iterations of the three methods for Rosenbrock example
Figure 2: Numerical performance of the runtime and number of iterations of the three methods for Rosenbrock example
@@ 5.1 Minimization of functional containing multiple summands
Now, consider the product manifold $\mathcal{M}^{N}=\{\textbf{x}=(x_{1},\ldots,x_{N})\colon x_{k}\in\mathcal{M},k\in\varLambda_{N}\}$. Define
$\textbf{D}=\{\textbf{x}\in\mathcal{M}^{N}\colon x_{1}=\dots=x_{N}\in\mathcal{M}\}.$ (53)
@@
$\operatorname*{Prox}_{\iota_{\textbf{D}}}(\textbf{x})=P_{\textbf{D}}(\textbf{x})=\Bigl(\operatorname*{argmin}_{x\in\mathcal{M}}\sum_{k\in\varLambda_{N}}d^{2}(x_{k},x),\ldots,\operatorname*{argmin}_{x\in\mathcal{M}}\sum_{k\in\varLambda_{N}}d^{2}(x_{k},x)\Bigl)\in\mathcal{M}^{N}.$ (54)
Define $F\colon\mathcal{M}^{N}\to(-\infty,\infty]$ by
$F(\textbf{x})=\sum_{k\in\varLambda_{N}}f_{k}(x_{k})\text{ for all }\textbf{x}=(x_{1},\ldots,x_{N})\in\mathcal{M}^{N}.$ (55)
We are interested in minimization of $F$ over $D$. For this, we consider the following unconstrained minimization problem:
$\min_{\textbf{x}\in\mathcal{M}^{N}}F(\textbf{x})+\iota_{\textbf{D}}(\textbf{x}).$ (56)
A detailed explanation of the above formulation is given in [4]. Note that for $\lambda>0$
$\operatorname*{Prox}_{\lambda F}(\textbf{x})$ $=\left(\operatorname*{Prox}_{\lambda f_{1}}(x_{1}),\ldots,\operatorname*{Prox}_{\lambda f_{N}}(x_{N})\right)\text{ for all }\textbf{x}\in\mathcal{M}^{N}.$
@@
$R_{\lambda F}(\textbf{x})=\left(R_{\lambda f_{1}}(x_{1}),\ldots,R_{\lambda f_{N}}(x_{N})\right)\text{ for all }\textbf{x}\in\mathcal{M}^{N}.$
@@
$\begin{cases}\textbf{y}_{n}=\exp_{\textbf{x}_{n}}(-\theta_{n}\exp_{\textbf{x}_{n}}^{-1}\textbf{x}_{n-1}),\\ \textbf{x}_{n+1}=\gamma(\textbf{y}_{n},{R}_{\lambda F}{R}_{\iota_{\textbf{D}}}(\textbf{y}_{n});\alpha_{n})\text{ for all }n\in\mathbb{N},\end{cases}$ (57)
where $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ are sequences satisfying the conditions (C1), (C2) and (C3). Then $\{\textbf{x}_{n}\}$ converges to an element $\textbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\iota_{\textbf{D}}})$ such that $\operatorname*{Prox}_{\iota_{\textbf{D}}}(\textbf{v})$ is solution of problem (56).
@@
Since $T=R_{\lambda F}R_{\iota_{\textbf{D}}}$ is nonexpansive. From Theorem 32 sequence $\{\textbf{x}_{n}\}$ generated by (57) converges to an element $\textbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\iota_{\textbf{D}}})$. From Theorem 16, we have $\textbf{u}=\operatorname*{Prox}_{\iota_{\textbf{D}}}(\textbf{v})$ is solution of problem (56). ∎
@@
$\begin{cases}\textbf{y}_{n}=\gamma(\textbf{x}_{n},{R}_{\lambda F}{R}_{\iota_{\textbf{D}}}(\textbf{x}_{n});\alpha_{n}),\\ \textbf{x}_{n+1}=({R}_{\lambda F}{R}_{\iota_{\textbf{D}}})^{p}(\textbf{y}_{n})\text{ for all }n\in\mathbb{N}_{0},\end{cases}$ (58)
where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ satisfying the condition $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$. Then $\{\textbf{x}_{n}\}$ converges to an element $\textbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\iota_{\textbf{D}}})$ such that $\operatorname*{Prox}_{\iota_{\textbf{D}}}(\textbf{v})$ is solution of problem (56).
@@
Since $T=R_{\lambda F}R_{\iota_{\textbf{D}}}$ is nonexpansive. From Theorem 33, sequence $\{\textbf{x}_{n}\}$ generated by (58) converges to an element $\textbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\iota_{\textbf{D}}})$. From Theorem 16, we have $\textbf{u}=\operatorname*{Prox}_{\iota_{\textbf{D}}}(\textbf{v})$ is solution of problem (56). ∎

### v2

@@ 3 Efficient Douglas-Rachford Algorithms on Hadamard Manifolds
Initialization: Choose $x_{1}\in\mathcal{M}$ arbitrary.
@@
$\begin{cases}y_{n}=\gamma(x_{n},{R}_{\lambda\phi}{R}_{\lambda\psi}(x_{n});\alpha_{n}),\\ x_{n+1}=({R}_{\lambda\phi}{R}_{\lambda\psi})^{p}(y_{n})\text{ for all }n\in\mathbb{N},\end{cases}$ (45)
@@
Remark 28.
If $\mathcal{M}=\mathbb{R}^{m}$, the Euclidean space, and we choose $A=\partial\phi$ and $B=\partial\psi$ in the inertial Douglas-Rachford algorithm of Boţ et al. [Bot2015, Theorem 8], then their method is equivalent to Algorithm 26. Note that we use parameter conditions different from those in Boţ et al. [Bot2015]; see Remark 17.
@@
Since $T={R}_{\lambda\phi}{R}_{\lambda\psi}$ is nonexpansive. From Theorem 18, sequence $\{x_{n}\}$ generated by (44) converges to an element $v\in\operatorname*{Fix}({R}_{\lambda\phi}{R}_{\lambda\psi})$. From Theorem 12, $u=\operatorname*{Prox}_{\lambda\psi}(v)$ is solution of problem (2). ∎
@@
Since $T={R}_{\lambda\phi}{R}_{\lambda\psi}$ is nonexpansive. From Theorem 23, sequence $\{x_{n}\}$ generated by (45) converges to an element $v\in\operatorname*{Fix}({R}_{\lambda\phi}{R}_{\lambda\psi})$. From Theorem 12, $u=\operatorname*{Prox}_{\lambda\psi}(v)$ is solution of problem (2). ∎
@@ 4 Numerical Experiments
where $a,b\in(0,\infty)$ (see [Rosenbrock]). In order to handle the nonconvex problem (46) in $\mathcal{M}=\mathbb{R}^{2}$, consider
@@
Then $\mathcal{M}$ is a Hadamard manifold of zero sectional curvature with respect to metric (47) (see [DaCruz]) and the following statements hold:
@@
Thus $\phi$ and $\psi$ are geodesically convex on $\mathcal{M}$. For more details of this splitting; see [BergmannJOTA2024].
@@
    For $\lambda>0$, the proximal operators $\operatorname*{Prox}_{\lambda\phi}$ and $\operatorname*{Prox}_{\lambda\psi}$ are defined by
    $\operatorname{Prox}_{\lambda\phi}(x)=\begin{pmatrix}x_{1}\\ \frac{x_{2}+2a\lambda x_{1}^{2}}{1+2a\lambda}\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}\text{ and }$
    $\operatorname{Prox}_{\lambda\psi}(x)=\begin{pmatrix}\frac{x_{1}+2\lambda b}{1+2\lambda}\\ x_{2}-\frac{4\lambda(x_{1}+2\lambda b)(x_{1}-b)+4\lambda^{2}(x_{1}-b)^{2}}{(1+2\lambda)^{2}}\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}.$
@@
    $R_{\lambda\phi}(x)=\begin{pmatrix}x_{1}\\ \frac{2(x_{2}+2a\lambda x_{1}^{2})}{1+2a\lambda}-x_{2}\end{pmatrix}\text{ for all }x=(x_{1},x_{2})^{T}\in\mathcal{M}\text{ and }$ (49)
@@
By Proposition 15, the reflections $R_{\lambda\phi}$ and $R_{\lambda\psi}$ defined by (49) and (50), respectively, are nonexpansive.
@@
$\min_{x\in\mathcal{M}}\phi(x)+\psi(x),$ (51)
where $\phi$ and $\psi$ are defined by (48).
For the numerical experiment, we compare three different methods: the Douglas-Rachford method [Alg(DR)] (3), the inertial Douglas-Rachford method [Alg(InDR)] (44) and the $p$-accelerated normal S-Douglas-Rachford method [Alg($p$-AccDR)] (45). All numerical tests are carried out using MATLAB R2024b on an Apple Macbook Air Laptop with a 1.8 GHz Dual-Core Intel Core i5 processor and 8 GB RAM. All algorithms stops when they meet the stopping criterion $E_{n}=d(s_{n},s_{n-1})<10^{-14}$, where $s_{n}=\operatorname*{Prox}_{\lambda\psi}(x_{n})$.
Numerical experiments are performed on the Rosenbrock function (46) with parameters $(a,b)=(1,2)$ and $(3,7)$. The initial point is chosen as $x_{0}=(1,2)^{\top}$ for all methods, while Method (InDR) uses $x_{1}=(1,3)^{\top}$ in both cases.
We set $\lambda=1$ for all three algorithms and iteration parameters are chosen as follows:
Algorithm $(a,b)=(1,2)$ $(a,b)=(1,2)$ $(a,b)=(3,7)$ $(a,b)=(3,7)$
Alg(DR) $\alpha_{n}=0.1$ $\alpha_{n}=0.3$ $\alpha_{n}=0.2$ $\alpha_{n}=0.4$
Alg(InDR) $\alpha_{n}=0.1,$ $\alpha_{n}=0.3,$ $\alpha_{n}=0.2,$ $\alpha_{n}=0.4,$
$\theta_{n}=0.42$ $\theta_{n}=0.35$ $\theta_{n}=0.4$ $\theta_{n}=0.33$
Alg($p$-AccDR) $\alpha_{n}=0.1$ $\alpha_{n}=0.3$ $\alpha_{n}=0.2$ $\alpha_{n}=0.4$
Note that in our parameter choice, we have taken the same $\alpha_{n}$ (in the Mann-type step) for all methods. Corresponding to this choice of $\alpha_{n}$, we selected the parameter $\theta_{n}$ according to condition (C3) for Alg(InDR), that is, $\epsilon=(1-\alpha_{n})/\alpha_{n}$, $\epsilon_{1}=1+\epsilon+\max\{1,\epsilon\}$, and $\theta_{n}\approx 0.9(\epsilon/\epsilon_{1})$. We chose $p=2$, which is the smallest value that yielded the best performance in our experiments. These parameters were chosen empirically by testing several reasonable values.
The runtime and number of iterations are presented in Table 1 for $(a,b)=(1,2)$ and in Table 2 for $(a,b)=(3,7)$. The graph of $E_{n}$ values with respect to the number of iterations $n$ is shown in Figure 2 for both cases. From Tables 1, 2, and Figure 2, it is evident that Alg($p$-AccDR) is faster than Alg(InDR) and Alg(DR) in terms of both runtime and number of iterations.
Table 1: Numerical performance of the runtime and number of iterations for Rosenbrock problem with $(a,b)=(1,2)$
$(\alpha_{n},\theta_{n})=(0.1,0.42)$ $(\alpha_{n},\theta_{n})=(0.3,0.35)$
Algorithm Runtime(sec) Iterations Runtime(sec) Iterations
Alg(DR) 0.000800 213 0.000468 63
Alg(InDR) 0.000461 102 0.000287 43
Alg(pp-AccDR) 0.000088 15 0.000099 13
Table 2: Numerical performance of the runtime and number of iterations for Rosenbrock problem with $(a,b)=(3,7)$
$(\alpha_{n},\theta_{n})=(0.2,0.4)$ $(\alpha_{n},\theta_{n})=(0.4,0.33)$
Algorithm Runtime(sec) Iterations Runtime(sec) Iterations
Alg(DR) 0.001003 106 0.000623 44
Alg(InDR) 0.000385 55 0.000368 37
Alg(pp-AccDR) 0.000150 30 0.000130 19
(a) $(a,b)=(1,2),(\alpha_{n},\theta_{n})=(0.1,0.42)$
(b) $(a,b)=(1,2),(\alpha_{n},\theta_{n})=(0.3,0.35)$
(c) $(a,b)=(3,7),(\alpha_{n},\theta_{n})=(0.2,0.4)$
(d) $(a,b)=(3,7),(\alpha_{n},\theta_{n})=(0.4,0.33)$
Figure 2: Comparison of $E_{n}$ with respect to number of iterations for the Rosenbrock Problem
@@ 5.1 Minimization of functional containing multiple summands
Now, consider the product manifold $\mathcal{M}^{N}=\{\mathbf{x}=(x_{1},\ldots,x_{N})\colon x_{k}\in\mathcal{M},k\in\varLambda_{N}\}$. Define
$\mathbf{D}=\{\mathbf{x}\in\mathcal{M}^{N}\colon x_{1}=\dots=x_{N}\in\mathcal{M}\}.$ (53)
Then $\mathbf{D}$ is a nonempty, closed and geodesically convex set. Hence, $\iota_{\mathbf{D}}$ is a proper, geodesically convex and lower semicontinuous function; see [BacakNAA2014]. For $\mathbf{x}=(x_{1},\ldots,x_{N})\in\mathcal{M}^{N}$
$\operatorname{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{x})=P_{\mathbf{D}}(\mathbf{x})=\Bigl(\operatorname*{argmin}_{x\in\mathcal{M}}\sum_{k\in\varLambda_{N}}d^{2}(x_{k},x),\ldots,\operatorname*{argmin}_{x\in\mathcal{M}}\sum_{k\in\varLambda_{N}}d^{2}(x_{k},x)\Bigl)\in\mathcal{M}^{N}.$
Define $F\colon\mathcal{M}^{N}\to(-\infty,\infty]$ by
@@
We are interested in minimization of $F$ over $\mathbf{D}$. For this, we consider the following unconstrained minimization problem:
$\min_{\mathbf{x}\in\mathcal{M}^{N}}F(\mathbf{x})+\iota_{\mathbf{D}}(\mathbf{x}).$ (55)
A detailed explanation of the above formulation is given in [Bergmann2016]. Note that for $\lambda>0$
$\operatorname{Prox}_{\lambda F}(\mathbf{x})$ $=\left(\operatorname*{Prox}_{\lambda f_{1}}(x_{1}),\ldots,\operatorname*{Prox}_{\lambda f_{N}}(x_{N})\right)\text{ for all }\mathbf{x}\in\mathcal{M}^{N}.$
@@
$R_{\lambda F}(\mathbf{x})=\left(R_{\lambda f_{1}}(x_{1}),\ldots,R_{\lambda f_{N}}(x_{N})\right)\text{ for all }\mathbf{x}\in\mathcal{M}^{N}.$ (56)
@@
$\begin{cases}\mathbf{y}_{n}=\exp_{\mathbf{x}_{n}}(-\theta_{n}\exp_{\mathbf{x}_{n}}^{-1}\mathbf{x}_{n-1}),\\ \mathbf{x}_{n+1}=\gamma(\mathbf{y}_{n},{R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}}(\mathbf{y}_{n});\alpha_{n})\text{ for all }n\in\mathbb{N},\end{cases}$ (57)
where $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ are sequences satisfying the conditions (C1), (C2) and (C3). Then $\{\mathbf{x}_{n}\}$ converges to an element $\mathbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}})$ such that $\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{v})$ is solution of problem (55).
@@
Since $T=R_{\lambda F}R_{\lambda\iota_{\mathbf{D}}}$ is nonexpansive. From Theorem 29, sequence $\{\mathbf{x}_{n}\}$ generated by (57) converges to an element $\mathbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}})$. From Theorem 12, we have $\mathbf{u}=\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{v})$ is solution of problem (55). ∎
@@
$\begin{cases}\mathbf{y}_{n}=\gamma(\mathbf{x}_{n},{R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}}(\mathbf{x}_{n});\alpha_{n}),\\ \mathbf{x}_{n+1}=({R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}})^{p}(\mathbf{y}_{n})\text{ for all }n\in\mathbb{N},\end{cases}$ (58)
where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ satisfying the condition $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$. Then $\{\mathbf{x}_{n}\}$ converges to an element $\mathbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}})$ such that $\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{v})$ is solution of problem (55).
@@
Since $T=R_{\lambda F}R_{\lambda\iota_{\mathbf{D}}}$ is nonexpansive. From Theorem 30, sequence $\{\mathbf{x}_{n}\}$ generated by (58) converges to an element $\mathbf{v}\in\operatorname*{Fix}({R}_{\lambda F}{R}_{\lambda\iota_{\mathbf{D}}})$. From Theorem 12, we have $\mathbf{u}=\operatorname*{Prox}_{\lambda\iota_{\mathbf{D}}}(\mathbf{v})$ is solution of problem (55). ∎